OpenStudy (anonymous):

The slope of a line is –2 and its y-intercept is (0, 3). What is the equation of the line that is parallel to the first line and passes through (2, 2)?

OpenStudy (ivettef365):

equation of the line is y = mx + b where m represents the slope and b the y-int.

OpenStudy (ivettef365):

so you have slope of -2, and the question ask for the equation parallel, that means slope remain as -2 and passes through (2,2) so you have equation y = mx + b you have points (2,2) and m=-2 only find b 2 = -2(2) + b